OpenStudy (anonymous):

f(x)=x^3 + x and h be the inverse of f. Calculate h'(10).

OpenStudy (anonymous):

I have no idea how to start these.

OpenStudy (anonymous):

Basically solve for \(x\) where \[ 10 = x^3+x \]

OpenStudy (anonymous):

i would make an educated guess first

OpenStudy (anonymous):

Actually, it's a derivative right?

OpenStudy (anonymous):

again you are going to use \[(f^{-1})'(x)=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

you need first to find \(f^{-1}(10)\) then you need \(f'(that)\) and finally flip it

OpenStudy (anonymous):

I get that it equals 301?

OpenStudy (anonymous):

but the problem is cooked up so that \(f^{-1}(10)\) should be real easy to find think of solving \(x^3+x=10\) in your head

OpenStudy (anonymous):

why do you set them equal to each other?

OpenStudy (anonymous):

slow not \(f(10)\) but rather \(f^{-1}(10)\) i.e. what would you put in for \(x\) to get \(f(x)=10\)

OpenStudy (anonymous):

I thought it was a derivative, I got it.

OpenStudy (anonymous):

You can get his formula by using chain rule: \[ f(h(x)) = x \implies [f(h(x)) ]' = 1 \implies f'(h(x))h'(x) = 1 \implies h'(x) = \frac{1}{f'(h(x))} \]

OpenStudy (anonymous):

okay, so I set them equal to each other. and solve for x?

OpenStudy (anonymous):

k good this is what you are using for all of these \[(f^{-1})'(x)=\frac{1}{f'(f^{-1}(x))}\] what you need is \[(f^{-1})'(10)=\frac{1}{f'(f^{-1}(10))}\] so the first number you have to find is the insidest one \[f^{-1}(10)\] which is another way of saying solve \(f(x)=10\) or in your case solve \(x^3+x=10\)

OpenStudy (anonymous):

solve this in your head then find \(f'(x)\) and then take \(f'(\text{what you found})\) then flip it

OpenStudy (anonymous):

2!

OpenStudy (anonymous):

eureka!

OpenStudy (anonymous):

now we are up to \[(f^{-1})'(10)=\frac{1}{f'(f^{-1}(10))}=\frac{1}{f'(2)}\]

OpenStudy (anonymous):

your next job is to find \(f'(2)\) which again you can probably do in your head

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

seems unlikely \[f(x)=x^3+x\] what is \[f'(x)\]?

OpenStudy (anonymous):

don't you set 2 equal to that?

OpenStudy (anonymous):

confused still lets go back to this \[(f^{-1})'(10)=\frac{1}{f'(f^{-1}(10))}=\frac{1}{f'(2)}\] the denominator is just \(f'(2)\) has nothing to do with the inverse or with solving just compute that number

OpenStudy (anonymous):

so 12? 1/12

OpenStudy (anonymous):

yay except i think it is 13

OpenStudy (anonymous):

yep sorry forgot to add 1

OpenStudy (anonymous):

lol hope this is getting clearer

OpenStudy (anonymous):

so the final answer is 13?

OpenStudy (anonymous):

\[(f^{-1})'(10)=\frac{1}{f'(f^{-1}(10))}=\frac{1}{f'(2)}=\frac{1}{3\times 2^2+1}=\frac{1}{13}\]

OpenStudy (anonymous):

wait, no it is 1/13

OpenStudy (anonymous):

one over f prime of f inverse of x

OpenStudy (anonymous):

I understand now, it will just take a bit of practice! Thank You so much for helping me! I really appreciate it!

OpenStudy (anonymous):

yw now you will get to \[\frac{d}{dx}\sin^{-1}(x)=\frac{1}{\cos(\sin^{-1}(x))}\]

OpenStudy (anonymous):

one over f prime of f inverse of x you will want to write this as \[\frac{d}{dx}\sin^{-1}(x)=\frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

Wow!

OpenStudy (anonymous):

that is if you recall that \(\cos(\sin^{-1}(x))=\sqrt{1-x^2}\)

OpenStudy (anonymous):

similarly \[\frac{d}{dx}\ln(x)=\frac{1}{e^{\ln(x)}}=\frac{1}{x}\]thinking of \(f(x)=e^x, f^{-1}(x)=\ln(x)\)

OpenStudy (anonymous):

Just had to do another problem like the on above and I knew how to do it! Thanks again!

OpenStudy (anonymous):

yw good work